OpenStudy (anonymous):

How to determine all values of p for which the integral is improper. integral dx/(x^2+2) -3 to p ?????????

OpenStudy (anonymous):

is the integral (1/(x^2 +2))dx ?

OpenStudy (anonymous):

true from -3 to p

OpenStudy (anonymous):

and im assuming as p approaches infinity?

OpenStudy (anonymous):

okay lets get on ahead with this you might be confused how do you even integrate this? or do you have any clue

OpenStudy (lgbasallote):

\[\int_{-3}^p \frac{dx}{x^2 +2}dx\]

OpenStudy (anonymous):

i assumed p approaches -infinite and +infinite~ i got -pi/2 and pi/2

OpenStudy (lgbasallote):

lol too many dx

OpenStudy (anonymous):

what was your integration?

OpenStudy (anonymous):

\[\lim_{t \rightarrow p}\]\[[\tan^{-1} (x/\sqrt{2})]_{-3}^{t}\]

OpenStudy (anonymous):

alright perfect good job first we will say as tan^(-1)(infinity) = pi/2 and tan^(-1)(- infinity) = -pi/2 so now let us see as p approaches infinity we have lim p --> inf = tan^(-1)( inf)/sqrt2 and therefore this is simply pi/2/sqrt 2 = pi/2sqrt2 and as lim p --> -inf = tan^(-1)( -inf)/sqrt2 and this is simply -pi/2/sqrt2 = -pi/2sqrt2 :) hope you followed

OpenStudy (anonymous):

thx buddy~ got it~